# Learning WENO for entropy stable schemes to solve conservation laws

### v1

@@ 3.3 Feasible Region for SP-WENO
Case $\bm{\Omega_{\Theta}}$ Sign Property Constraint Accuracy Constraint Remarks
@@ 1 $\theta_{i}^{+},\theta_{i+1}^{-}>1$ $C_{1}=C_{2}=\frac{1}{8}$ None $\tilde{w}_{0}=w_{1}=0$
Table 1: Summary of the possible $\Omega_{\Theta}$ along with the sign property and accuracy constraints defining the associated feasible region $\Omega_{C}$. Here $\mathcal{L}$ is given by (26).
4 DSP-WENO
To this end, an artificial neural network ${\bm{\mathcal{N}}}_{{\bm{\Phi}}}:\mathbb{R}^{N_{I}}\mapsto\mathbb{R}^{N_{O}}$, consisting of multiple affine transformations (layers) separated by non-linear activation functions, is trained to approximate the unknown function. Each layer comprises a set of neurons that are parameterized by tunable weights and biases. Mathematically, the simplest neural network architecture, known as the multilayer perceptron (MLP), is of the form
${\bm{\mathcal{N}}}_{{\bm{\Phi}}}:=\rho\circ\mathcal{A}_{\Phi_{L+1}}^{(L+1)}\circ\sigma\circ\mathcal{A}_{\Phi_{L}}^{(L)}\circ\sigma\circ\mathcal{A}_{\Phi_{L-1}}^{(L-1)}\circ\cdots\circ\sigma\circ\mathcal{A}_{\Phi_{1}}^{(1)},$ (28)
where $\mathcal{A}_{\Phi_{k}}^{(k)}$ is an affine transformation representing a layer of the MLP parameterized by its weights and biases $\Phi_{k}$, $\sigma$ is a non-linear activation function applied component-wise, and $\rho$ is an output function that transforms the network output into a usable form. The first $L$ parameterized layers are known as the hidden layers, while the final layer is known as the output layer. $\bm{\Phi}=\{\Phi_{k}\}_{k=1}^{L}$ is the collection of all the trainable weights and biases of the network.
Given a training dataset, the network parameters are learned by optimizing an objective/loss function $\Pi(\bm{\Phi})$ via an iterative gradient descent-type optimization algorithm. These algorithms require the gradient of the MLP with respect to its trainable parameters, which can be efficiently computed using back-propagation. Assuming that a suitable network architecture is chosen (along with any other hyperparameters), and with sufficient representative data in the training set, the trained network is capable of approximating the true function ${\bm{\mathcal{N}}}$ [42, 24, 57].
For the $k$-th sample in $\mathbb{T}$, $\bm{X}^{(k)}\in{\mathbb{R}}^{4}$ consists of the four cell center values of some function $z$ in the local stencil centered about (some) $x_{i+\frac{1}{2}}$: $[z_{i-1},z_{i},z_{i+1},z_{i+2}]$, while $\bm{Y}^{(k)}\in{\mathbb{R}}^{2}$ is the true cell interface values $z^{\pm}(x_{i+\frac{1}{2}})$. Note that the interface values from the left (-) and right (+) will be identical for a smooth function, but might be different if there is a discontinuity at the interface.
We want to ensure that all the inputs to the neural network are of the same order of magnitude (essential for generalization), so we first scale the cell center values as
Note that a similar input scaling is also considered in [45, 46]. Using the scaled cell center values $[z_{i-1}^{*},z_{i}^{*},z_{i+1}^{*},z_{i+2}^{*}]$, we compute $\theta_{i+1}^{-},\theta_{i}^{+},\left|\Delta z_{i-\frac{1}{2}}^{*}\right|,\left|\Delta z_{i+\frac{1}{2}}^{*}\right|,\left|\Delta z_{i+\frac{3}{2}}^{*}\right|$. We note that the jump ratios $\theta_{i+1}^{-},\theta_{i}^{+}$ are invariant to the scaling (29). Thus we define the map ${\bm{\mathcal{S}}}:{\mathbb{R}}^{4}\rightarrow{\mathbb{R}}^{5}$ which transforms $\bm{X}^{(k)}$ in the training set to the jump ratios and scaled absolute jumps mentioned above, i.e.,
[…]
The output of the neural network ${\bm{\mathcal{N}}}_{\bm{\Phi}}$ is the vector of convex weights $\bm{\alpha}\in{\mathbb{R}}^{5}$. In other words, we have
[…]
Let us also define a function ${\bm{\mathcal{V}}}:{\mathbb{R}}^{5}\rightarrow{\mathbb{R}}^{2\times 5}$ that maps ${\bm{\mathcal{S}}}(\bm{X}^{(k)})$ to 5 vertices in ${\mathbb{R}}^{2}$, i.e.,
In (33), the $\bm{\nu}^{(k)}_{s}$ corresponds to the vertices of a convex polygon (defining the feasible region), thus ensuring that any convex combination of these vertices will result in a vector in ${\mathbb{R}}^{2}$ that lies in the (closure) of this convex polygon. These convex regions are formed by at most five vertices. In situations with less than five vertices, some of the $\bm{\nu}^{(k)}_{s}$ in (33) are replaced by a repeated interior node of the convex region, typically the centroid. We provide more precise details of ${\bm{\mathcal{V}}}$ in Section 4.2. The output of the MLP is combined with these vertices to obtain the DSP-WENO weight perturbations $C_{1},C_{2}$, given by
[…]
We finally define the objective function in terms of the trainable parameters ${\bm{\Phi}}$ of the network
$\Pi(\bm{\Phi})$ $=\frac{1}{K}\sum_{k=1}^{K}\Pi_{k}({\bm{\Phi}}),$ (36)
$\Pi_{k}({\bm{\Phi}})$ $=\left\|\widehat{\bm{Y}}^{(k)}-\bm{Y}^{(k)}\right\|_{1}$
[…]
where .$\|.\|_{1}$ denotes the $l_{1}$ norm. Thus, the network is trained by solving the minimization problem
${\bm{\Phi}}^{\dagger}=\underset{{\bm{\Phi}}}{\arg\min}\Pi(\bm{\Phi}).$
Remark 2.
We have tried the training process with other loss functions, including the mean squared error loss function, but ultimately using the loss function (36) provides the best performance. Moreover, this is particularly meaningful since the $L^{1}$ norm is typically used when computing convergence in conservation laws.
Remark 3.
Figure 2: Schematic for the DSP-WENO reconstruction algorithm.
4.1 Data Selection
To generate training data for the neural network, we sample stencils consisting of four cells from parameterized smooth and discontinuous functions of the form listed in Table 2. The dataset has a 50/50 balanced split of smooth and discontinuous data. The section of the dataset consisting of smooth data consists of an equal representation of the three smooth types listed in Table 2. For discontinuous data, the jump will either be between the first and second cells, the second and third cells, or the third and fourth cells. In any case, using the four cell center values, we compute the jump ratios $\theta_{i+1}^{-}$ and $\theta_{i}^{+}$ and the absolute jumps across the three cell interfaces in the stencil, via ${\bm{\mathcal{S}}}$ defined in (30). Further, we compute a set of vertices using ${\bm{\mathcal{V}}}$ defined in (33). Finally, the true values at the cell interface $x_{i+\frac{1}{2}}$ are recorded to serve as target values that DSP-WENO will aim to reconstruct. Overall, the dataset consists of 100,000 samples, roughly corresponding to 50,000 discontinuous samples and 50,000 smooth samples.
No. Type $u(x)$ Parameters
1 Smooth $ax^{3}+bx^{2}+cx+d$ $a,b,c,d\in\mathbb{U}[-10,10]$
2 Smooth $(x-a)(x-b)(x-c)+d$ $a,b,c,d\in\mathbb{U}[-2,2]$
3 Smooth $\sin{(a\pi x+b)}$ $a,b\in\mathbb{U}[-2,2]$
Remark 4.
We do not use solutions to conservation laws to train the network. We instead generate functions of varying regularity that canonically represent the local solution features we can expect to observe in solutions to conservation laws. Thus, there is an insignificant cost in generating the training data. A similar strategy was also considered in [45, 46].
Remark 5.
In our experiments, we noticed that the performance of DSP-WENO, particularly in the framework of TeCNO schemes, was highly dependent on the training set used to train DSP-WENO. In other words, the training set is itself a hyperparameter, that needs to be carefully constructed based on domain expertise.
4.2 Vertex Selection Algorithm
In this section, we describe the vertex selection algorithm that defines the map ${\bm{\mathcal{V}}}$ in (33). As discussed in Section 3.3 and listed in Table 1, there is a feasible region $\Omega_{C}$ associated with each $\Omega_{\Theta}$. Any choice of $(C_{1},C_{2})$ in the feasible region leads to consistent WENO weights which are guaranteed to satisfy the sign property and yield third-order accuracy in smooth regions. We show below that the $\Omega_{C}$ constructed to satisfy the various feasibility constraints leads to a convex polygon in $\mathbb{R}^{2}$. The objective with the neural network approach is to learn how to explore these $\Omega_{C}$ to better select $(C_{1},C_{2})$. In comparison, SP-WENO and SP-WENOc have handcrafted weight selection strategies that suffer from significant overshoots near discontinuities in the TeCNO framework. DSP-WENO aspires to mitigate this behavior through a data-driven search in the feasible regions.
@@ 4.2 Vertex Selection Algorithm
Now, we discuss how the vertices of the convex $\Omega_{C}$ are selected. The fully-detailed algorithm can be found in Algorithm 1. Note that the vertex algorithm is applied for each four-cell stencil in the mesh. With the exception of cases 2 and 3 (see Table 1), the vertex selection for most scenarios is quite straight-forward. This is due to the fact that all cases besides 2 and 3 do not require an additional accuracy constraint, and so we only need to select vertices that yield a convex figure that satisfies the sign property and consistency of the WENO weights. Cases 2 and 3 are more involved since we do require that $C_{1},C_{2}=\mathcal{O}(h)$ in smooth regions. To address this, we construct an $\mathcal{O}(h)$ region in $\mathbb{R}^{2}$ as depicted in Figure 3. The dashed lines enclose a box representing the consistency requirements for the perturbations, i.e., $-3/8\leq C_{1},C_{2}\leq 1/8$. The $\mathcal{L}=1$ line shown in orange separates the regions satisfying the sign property for Cases 2 and 3, represented as solid dark and light grey regions respectively. Within these grey regions, we construct smaller sub-regions that also satisfy the accuracy constraints. These sub-regions are shown in blue and green, and form the required feasible regions $\Omega_{C}$ for cases 2 and 3. To construct these sub-regions, we first define the box $\mathcal{H}=[\gamma_{2},\gamma_{1}]\times[\gamma_{2},\gamma_{1}]\subset[-3/8,1/8]\times[-3/8,1/8]$ where
$\gamma_{1}$ $=\min{\left(\max{\left(|\Delta z_{i-\frac{1}{2}}^{*}|,|\Delta z_{i+\frac{1}{2}}^{*}|,|\Delta z_{i+\frac{3}{2}}^{*}|\right)},\frac{1}{8}\right)},$ (38)
@@
The region $\mathcal{H}$ is shown as a red dotted box in Figure 3. Note that when the $z$ on the local stencil is smooth, $|\Delta z_{i-\frac{1}{2}}^{*}|,|\Delta z_{i+\frac{1}{2}}^{*}|,|\Delta z_{i+\frac{3}{2}}^{*}|=\mathcal{O}(h)$. Thus, we ensure that $\gamma_{1},\gamma_{2}=\mathcal{O}(h)$, and searching for $(C_{1},C_{2})$ inside $\mathcal{H}$ would constrain $C_{1},C_{2}=\mathcal{O}(h)$, which is the order constraint for cases 2 and 3 assuming smoothness for $z$. When a discontinuity is present, $\mathcal{H}$ can potentially expand to the whole consistency region for $C_{1},C_{2}$. With the line $\mathcal{L}=1$ splitting $\mathcal{H}$, we have that the feasible region for Case 2 consists of the region above $\mathcal{L}=1$ and that of Case 3 consists of the region below $\mathcal{L}=1$. We further note that the closer the perturbations are to zero, the more accurate the reconstruction. In smooth regions, this is desirable, but near discontinuities, capturing shocks becomes the priority, so searching a wider search space becomes important.
@@
As shown in Figure 3, the line $\mathcal{L}=1$ splits $\mathcal{H}$ into a triangular and pentagonal $\Omega_{C}$ region, with two out of three and two out of five vertices lying on the line $\mathcal{L}=1$, respectively. A peculiarity of choosing $(C_{1},C_{2})$ on this line is that it ensures the left and right reconstructed states at the interface are equal, i.e, $\llbracket z\rrbracket_{i+\frac{1}{2}}=0$. As discussed in Section 3.2, a zero reconstruction jump is undesirable in the vicinity of a discontinuity in the TeCNO framework. When choosing $(C_{1},C_{2})$ from a triangular region, two of whose vertices are on this line, which biases the selection to be on this line (this is what we also observed in practice). Thus, when constructing a triangular $\Omega_{C}$ above or below $\mathcal{L}=1$, we do not use the triangle that is a part of $\mathcal{H}$. Instead, we replace one of the vertices lying on $\mathcal{L}=1$ with another vertex pulled away from this line (without violating the consistency and sign property constraints) to form a flipped triangular $\Omega_{C}$ that is outside of $\mathcal{H}$. This flipped $\Omega_{C}$ still satisfies $C_{1},C_{2}=\mathcal{O}(h)$, and thus any point selected in this triangle will also satisfy this constraint. For instance, let us consider Case 2 when $\psi^{+}<-1$, where $\Omega_{C}$ is triangular (see Figure 3(c),(d)). In this scenario, according to Algorithm 1, the vertex outside $\mathcal{H}$ is $(\gamma_{2},\frac{1}{8}(1+\psi^{-})-\gamma_{1}\psi^{-})$. For smooth regions, $C_{1}=\gamma_{2}=\mathcal{O}(h)$. Further, we can deduce that $(1+\psi^{-})=\mathcal{O}(h)$. Since $\gamma_{1}=\mathcal{O}(h)$, it is clear that $C_{2}=\frac{1}{8}(1+\psi^{-})-\gamma_{1}\psi^{-}=\mathcal{O}(h)$. Similar arguments can be made for Case 3 when $-1<\psi^{+}<0$ where $\Omega_{C}$ is once again triangular (see Figure 3(a),(b)).
The number of vertices needed to construct a convex figure across all the possible scenarios is at most five. Thus, to ensure that the neural network always outputs the same number of convex weights $\bm{\alpha}$ corresponding to the vertices, we constrain the output dimension of the network to be five. In the various scenarios where the number of vertices is less than five, we augment the set of vertices to a set of five vertices by including either redundant vertices or some averaged state of the figure’s vertices. For instance, in Cases 2 and 3, when the feasible region is a triangle, we augment the set of three vertices with the centroid of the triangle twice.
Remark 6.
@@
Remark 8.
The vertex modification algorithm for the triangular $\Omega_{C}$ is not unique. After experimenting with a few configurations of choosing the shifted vertex, we found the one used in Algorithm 1 yields the best performance.
Figure 3: Feasible region for Case 2 (dark grey) and Case 3 (light grey). The $\mathcal{O}(h)$ box $\mathcal{H}$ (outlined by red dots) is centered about the origin. The vertex selection algorithm further restricts the search space to the blue and green regions for Case 2 and Case 3, respectively. (a) $\psi^{+}\in(-1,0)$ with a zoomed in view in (b), (c) $\psi^{+}\in(-\infty,-1)$ with a zoomed in view in (d).
Algorithm 1 Vertex Selection Algorithm
@@
1:Evaluate $\psi^{\pm}$ according to (16), and $\gamma_{1},\gamma_{2}$ according to (38)

### v2

@@ Acknowledgements
Case $\bm{\Omega_{\Theta}}$ Sign Prop. Const. Acc. Const. Remarks
@@ 1 $\theta_{i}^{+},\theta_{i+1}^{-}>1$ $C_{1}=C_{2}=\frac{1}{8}$ None $\tilde{w}_{0}=w_{1}=0$
Table SM1: Summary of the possible $\Omega_{\Theta}$ along with the sign property and accuracy constraints defining the associated feasible region $\Omega_{C}$. Here $\mathcal{L}$ is given by (SM1).
SM2 Description of the Vertex Selection Algorithm
We provide a complete and in-depth description of how the vertices of the convex $\Omega_{C}$ are selected. The pseudocode for the algorithm can be found in Algorithm SM1. Note that the vertex algorithm is applied for each four-cell stencil in the mesh. Further, the number of vertices needed to construct convex polygons described below across all the possible scenarios is at most five. Thus, to ensure that the neural network always outputs the same number of convex weights $\bm{\alpha}$ corresponding to the vertices, we constrain the output dimension of the network to be five. In the various scenarios where the number of vertices of $\Omega_{C}$ is less than five, we augment the set of vertices to a set of five vertices by including either redundant vertices or some averaged state of the figure’s vertices.
With the exception of cases 2 and 3 (see Table SM1), the vertex selection for most scenarios is quite straight-forward. This is due to the fact that all cases besides 2 and 3 do not have an additional accuracy constraint. In particular, for
  • Case 1: The feasible region $\Omega_{C}$ reduces to a single node $\left(\frac{1}{8},\frac{1}{8}\right)$. Thus, all five vertices are taken as this node.
  • Case 4a: $\Omega_{C}$ reduces the line segment $C_{1}=\frac{1}{8},-\frac{3}{8}\leq C_{2}\leq\frac{1}{8}$. Thus we take two vertices to be the right end $\left(\frac{1}{8},\frac{1}{8}\right)$, two vertices to be the left end $\left(-\frac{3}{8},\frac{1}{8}\right)$, and the final vertex to be the line segment’s mid-point $\left(-\frac{1}{8},\frac{1}{8}\right)$.
  • Case 5a: Is identical to Case 4a, but with the role of $C_{1}$ and $C_{2}$ interchanged.
  • Case 4b, 5b and 6: There are not constraints (except for consistency) and $\Omega_{C}=\left[-\frac{3}{8},\frac{1}{8}\right]\times\left[-\frac{3}{8},\frac{1}{8}\right]$. Thus four vertices are chosen as the four corners of this square region, while the fifth vertex is taken to be the centroid of this square.
@@
Cases 2 and 3 are more involved since we require that $C_{1},C_{2}=\mathcal{O}(h)$ in smooth regions. To address this, we construct an $\mathcal{O}(h)$ region in $\mathbb{R}^{2}$ as depicted in Figure SM1. The black dashed lines enclose a box representing the consistency requirements for the perturbations, i.e., $-3/8\leq C_{1},C_{2}\leq 1/8$. The $\mathcal{L}=1$ line shown in orange separates the regions satisfying the sign property for Cases 2 and 3, represented as solid dark and light grey regions respectively. Within these grey regions, we construct smaller sub-regions that also satisfy the accuracy constraints. These sub-regions are shown in blue and green, and form the required feasible regions $\Omega_{C}$ for cases 2 and 3. To construct these sub-regions, we first define the box $\mathcal{H}=[\gamma_{2},\gamma_{1}]\times[\gamma_{2},\gamma_{1}]\subset[-3/8,1/8]\times[-3/8,1/8]$ where
$\gamma_{1}$ $=\min{\left(\max{\left(|\Delta z_{i-\frac{1}{2}}^{*}|,|\Delta z_{i+\frac{1}{2}}^{*}|,|\Delta z_{i+\frac{3}{2}}^{*}|\right)},\frac{1}{8}\right)},$ (SM2)
@@
The region $\mathcal{H}$ is shown as a red dotted box in Figure SM1. Note that when the $z$ on the local stencil is smooth, $|\Delta z_{i-\frac{1}{2}}^{*}|,|\Delta z_{i+\frac{1}{2}}^{*}|,|\Delta z_{i+\frac{3}{2}}^{*}|=\mathcal{O}(h)$. Thus, we ensure that $\gamma_{1},\gamma_{2}=\mathcal{O}(h)$, and searching for $(C_{1},C_{2})$ inside $\mathcal{H}$ would constrain $C_{1},C_{2}=\mathcal{O}(h)$, which is the order constraint for cases 2 and 3 assuming smoothness for $z$. When a discontinuity is present, $\mathcal{H}$ can potentially expand to the whole consistency region for $C_{1},C_{2}$. With the line $\mathcal{L}=1$ splitting $\mathcal{H}$, we have that the feasible region for Case 2 consists of the region above $\mathcal{L}=1$ and that of Case 3 consists of the region below $\mathcal{L}=1$. We further note that the closer the perturbations are to zero, the more accurate the reconstruction. In smooth regions, this is desirable, but near discontinuities, capturing shocks becomes the priority, so searching a wider search space becomes important.
@@
As shown in Figure SM1, the line $\mathcal{L}=1$ splits $\mathcal{H}$ into a triangular and pentagonal $\Omega_{C}$ region, with two out of three and two out of five vertices lying on the line $\mathcal{L}=1$, respectively. A peculiarity of choosing $(C_{1},C_{2})$ on this line is that it ensures the left and right reconstructed states at the interface are equal, i.e, $\llbracket z\rrbracket_{i+\frac{1}{2}}=0$. As discussed in Section 3.2 in the main text, a zero reconstruction jump is undesirable in the vicinity of a discontinuity in the TeCNO framework. When choosing $(C_{1},C_{2})$ from a triangular region, two of whose vertices are on this line, which biases the selection to be on this line (this is what we also observed in practice). Thus, when constructing a triangular $\Omega_{C}$ above or below $\mathcal{L}=1$, we do not use the triangle that is a part of $\mathcal{H}$. Instead, we replace one of the vertices lying on $\mathcal{L}=1$ with another vertex pulled away from this line (without violating the consistency and sign property constraints) to form a flipped triangular $\Omega_{C}$ that is outside of $\mathcal{H}$. This flipped $\Omega_{C}$ still satisfies $C_{1},C_{2}=\mathcal{O}(h)$, and thus any point selected in this triangle will also satisfy this constraint. For instance, let us consider Case 2 when $\psi^{+}<-1$, where $\Omega_{C}$ is triangular (see Figure SM1(c),(d)). In this scenario, according to Algorithm SM1, the vertex outside $\mathcal{H}$ is $(\gamma_{2},\frac{1}{8}(1+\psi^{-})-\gamma_{1}\psi^{-})$. For smooth regions, $C_{1}=\gamma_{2}=\mathcal{O}(h)$. Further, we can deduce that $(1+\psi^{-})=\mathcal{O}(h)$. Since $\gamma_{1}=\mathcal{O}(h)$, it is clear that $C_{2}=\frac{1}{8}(1+\psi^{-})-\gamma_{1}\psi^{-}=\mathcal{O}(h)$. Similar arguments can be made for Case 3 when $-1<\psi^{+}<0$ where $\Omega_{C}$ is once again triangular (see Figure SM1(a),(b)). To ensure that we are always working with five vertices, when the feasible region is a triangle, we augment the set of three vertices with the centroid of the triangle twice.
Remark SM2.
@@
Remark SM4.
The vertex modification algorithm for the triangular $\Omega_{C}$ is not unique. After experimenting with a few configurations of choosing the shifted vertex, we found the one used in Algorithm SM1 yields the best performance.
Figure SM1: Feasible region for Case 2 (dark grey) and Case 3 (light grey). The $\mathcal{O}(h)$ box $\mathcal{H}$ (outlined by red dots) is centered about the origin. The vertex selection algorithm further restricts the search space to the blue and green regions for Case 2 and Case 3, respectively. (a) $\psi^{+}\in(-1,0)$ with a zoomed in view in (b), (c) $\psi^{+}\in(-\infty,-1)$ with a zoomed in view in (d).
Algorithm SM1 Vertex Selection Algorithm
@@
1:Evaluate $\psi^{\pm}$ according to (14) in the main text, and $\gamma_{1},\gamma_{2}$ according to (SM2)